# Generation of Real-time Robotic Emotional Expressions Learning from Human Demonstration in Mixed Reality

### v1

@@ Abstract
Expressive behaviors in robots are critical for effectively conveying their emotional states during interactions with humans. In this work, we present a framework that autonomously generates realistic and diverse robotic emotional expressions based on expert human demonstrations captured in Mixed Reality (MR). Our system enables experts to teleoperate a virtual robot from a first-person perspective, capturing their facial expressions, head movements, and upper-body gestures, and mapping these behaviors onto corresponding robotic components including eyes, ears, neck, and arms. Leveraging a flow-matching-based generative process, our model learns to produce coherent and varied behaviors in real-time in response to moving objects, conditioned explicitly on given emotional states. A preliminary test validated the effectiveness of our approach for generating autonomous expressions.
@@ II-A4 WebSocket Back-End
II-B Flow-Matching Policy
@@
II-B1 Flow Matching Model
@@
II-B2 Emotional-Expression Dataset

### v2

@@ Abstract
Expressive behaviors in robots are critical for effectively conveying their emotional states during interactions with humans. In this work, we present a framework that autonomously generates realistic and diverse robotic emotional expressions based on expert human demonstrations captured in Mixed Reality (MR). Our system enables experts to teleoperate a virtual robot from a first-person perspective, capturing their facial expressions, head movements, and upper-body gestures, and mapping these behaviors onto corresponding robotic components including eyes, ears, neck, and arms. Leveraging a flow-matching-based generative process, our model learns to produce coherent and varied behaviors in real-time in response to moving objects, conditioned explicitly on given emotional states. A preliminary test validated the effectiveness of our approach for generating autonomous expressions. Supplementary material can be found at https://wallacewangchao.github.io/fm-expressions/
@@ II-A4 WebSocket Back-End
II-B Hand Mapping
II-C Real Robot Implementation
Figure 3: Top: The MR headset subscribes to the actual rotation of the robot head and computes the rotation difference with respect to the human head. Bottom: The teleoperation view locally compensates the virtual screen orientation according to this difference, aligning the displayed video stream with the robot head pose and reducing visual mismatch to mitigate motion sickness.
Our approach can be directly integrated into a real robotic system for data collection and teleoperation. The operator’s head pose and facial expression data are transmitted to the robot using the same mapping strategy described in the previous sections.
The visual feedback is captured by a camera mounted on the robot’s neck and streamed to the operator through a WebRTC-based communication pipeline. However, when deploying the system on a real robot, a critical challenge arises from the discrepancy between human head motion and robot neck dynamics. Specifically, the robot neck typically responds more slowly than human head rotations.
If the virtual screen rigidly follows the human head pose while the robot head lags behind, the displayed visual content no longer corresponds to the robot’s actual viewing direction. This inconsistency can lead to a strong visual–vestibular mismatch and significantly increase the risk of motion sickness for the operator.
To address this issue, the MR application continuously subscribes to the robot’s actual head rotation and computes its difference relative to the human head orientation. Instead of fixing the virtual screen directly in front of the operator or strictly following the human head motion, the screen orientation is adjusted based on this rotation difference. As a result, the virtual screen rotates more slowly than the human head, remaining aligned with the robot’s real viewing direction.
This adaptive compensation strategy not only reduces visual inconsistency and motion sickness, but also allows the operator to better perceive and anticipate the robot’s physical head movements during teleoperation.
II-D Flow-Matching Policy
@@
II-D1 Flow Matching Model
@@
II-D2 Emotional-Expression Dataset